OpenStudy (anonymous):

HELP! WILL MEDAL! Let g be a function that is defined for all x, x ≠ 2, such that g(3) = 4 and the derivative of g is g′(x)=x^2-16/x-2 , with x ≠ 2. Find all values of x where the graph of g has a critical value.

OpenStudy (anonymous):

So the question is asking for the x values when g'(x) = 0?

OpenStudy (anonymous):

I think it just want the critical points, but I need also to figure out the original formula.

OpenStudy (anonymous):

The critical points are when the derivative equals 0

OpenStudy (anonymous):

I found the critical points to be 4 and -4. I forgot to put the rest of the questions to the problem on here. 2. For each critical value, state whether the graph of g has a local maximum, local minimum or neither. You must justify your answers with a complete sentence. 3. On what intervals is the graph of g concave down? Justify your answer. 4. Write an equation for the tangent line to the graph of g at the point where x = 3. 5. Does this tangent line lie above or below the graph at this point? Justify your answer

OpenStudy (anonymous):

Aha, so doing these steps will result in the original function ultimately

OpenStudy (anonymous):

For point 2, you need to check if it's going up or down on both sides of the critical points

OpenStudy (anonymous):

How do I do that @aleroth

OpenStudy (mathmate):

Hints: 1. how to find g(x) g'(x)=(x^2-16)/(x-2) g(x)=\(\int g'(x)dx\) and the constant of integration can be found using the relation g(3)=4 2. Integration To help integrate, break g'(x) into partial fractions: \(g'(x)=(x^2-16)/(x-2)=x-\frac{12}{x-2}\)+2 Once g(x) is found, the remaining questions are straightforward. On examining g(x), the statement appears flawed if g(x) is real: "Let g be a function that is defined for \(all\) x, x ≠ 2,... " 3. critical points are relevant only when g(x) is defined, i.e. when x \(\in\) dom g(x). Critical \(values\) are the values of the function g(x) at the critical points. So finding g(x) is important.

OpenStudy (anonymous):

@mathmate so is the x-12/x-2+2 my g(x).

OpenStudy (mathmate):

g'(x)=\(\dfrac{x2−16}{x−2}=x−\dfrac{12}{x−2}+2\) We need to integrate term by term to get: \(\int g'(x)dx =\int (x−\frac{12}{x−2}+2)dx = \dfrac{x^2+4x}{2}+\int \frac{12}{x-2}\) \(= \dfrac{x^2+4x}{2}+12~\log(x-2)+C\) which means that dom g(x) = \((2,+\infty \)) Note: I write log to mean \(\log_e\), and \(\log_{10}\) for log to base 10. The rest should follow with the hints above. If you need more help, just post.

OpenStudy (anonymous):

@mathmate I am still sort of confused by this problem. Can we just start from the begining?

OpenStudy (freckles):

To find the critical numbers just solve g'(x)=0 and solve g'(x) dne . The x values of these equations that are in the domain of the original function are the critical numbers.

OpenStudy (freckles):

And actually it says x=2 is not in the domain of the original function. So you really only need to solve g'(x)=0

OpenStudy (freckles):

that is solve x^2-16=0 and you are done these will be the x values where you have critical values

OpenStudy (freckles):

your question does say to find the critical values

OpenStudy (freckles):

so we don't need to find the original function

OpenStudy (mathmate):

@freckles I believe we need to find g(x) in order to know its domain, here's why. 1. Solution of g'(x)=0 gives x=\(\pm4\), while g(x) has a domain of \(x\in(2,\infty)\), which means that x=-4 is not a critical point and must be eliminated with a knowledge of g(x). 2. The critical value g(4) is required as part of the answer, so g(x) is required to find that. ... unless there is a different way to do that. @krissy3039 If all this is confusing to you, we can start over step by step, when we have a time both online. Feel free to tag me when you get online.

OpenStudy (freckles):

@mathmate I don't see it saying anything about finding critical values. "Let g be a function that is defined for all x, x ≠ 2, such that g(3) = 4 and the derivative of g is g′(x)=x^2-16/x-2 , with x ≠ 2. Find all values of x where the graph of g has a critical value."

OpenStudy (freckles):

also x has to be greater than 2?

OpenStudy (freckles):

I thought the domain was (-inf,2) union (2,inf)

OpenStudy (freckles):

oh and i didn't see those other questions

OpenStudy (mathmate):

"Find all values of x where the graph of g has a critical value." You're quite right! I misread the question. We just need the values of x, not the critical-values, mea culpa! However, we still need g(x) to find the domain that contains the critical points.

OpenStudy (mathmate):

As I mentioned earlier, I do admit that the question is misleading in saying that "Let g be a function that is defined for all x, x ≠ 2,...". You may have included -4 as a critical point because of the statement. This statement contradicts with the actual function g(x).

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{x} dx=\ln|x|+C\]

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{x-2}=\ln|x-2|+C\]

OpenStudy (mathmate):

@krissy3039 When you get online, you can tag me or @freckles . We're on the same wavelength, so no worries.

OpenStudy (freckles):